# A Survey of Vibe Coding with Large Language Models

### v1

@@ 4.2.1 Pre-training Objectives
LeetCodeDataset [306] 2025; 2,613 training problems and 256 test problems Python LeetCode algorithmic problems Wrapped LeetCode problems into data structures containing imports, entry_point, tests, completion, examples, and metadata Each problem provides complete unit tests and target code; suitable for model training and evaluation Function-level; includes Python test scripts
@@
OpenCodeReasoning [307] 2024; 735,255 samples covering 28,319 programming problems Multi-language programming problems with reasoning Integrated data from 11 programming platforms and generated solutions using NVIDIA R1 Large-scale high-quality programming reasoning data; supports training models with reasoning capabilities Problem-level; includes problem text, solutions, and reasoning
@@
DeepCoder [308] 2025; 24k verified coding problems, each with $\geq$5 test cases Competitive programming problems Merged TACO verified, Prime Intellect synthetic problems, LiveCodeBench, etc., filtered to 24k problems with complete tests Provides fully open high-quality data for RL training; reward uses binary model (pass/fail) Problem-level; includes complete test sets
@@
SWE-Bench Variants [11, 309, 310] Since 2023; Full: 2,294, Lite: 300, Verified: 500, Multimodal: 617, Multilingual: 300 Primarily Python, also multi-language versions; project-level issue resolution Collected issues and PRs from GitHub, recording base commits, patches, and test patches; some instances expert-verified Instances include problem descriptions, base commits, patches, tests, and metadata; Multilingual version extends to 9 languages Project-level; used for evaluating bug-fixing capabilities of coding agents
@@
Multi-SWE-Bench [311] 2025; 1,632 high-quality instances Seven languages: Java, TypeScript, JavaScript, Go, Rust, C, C++ Five-stage process: repository selection, PR crawling, environment determination, PR filtering, and human validation Fills the Python-only gap in SWE-Bench; cross-language evaluation for coding agents Project-level; used for evaluation
@@
MLE-bench [312] 2024; 75 Kaggle competitions, 3.3TB dataset Multi-language; ML engineering tasks including model training, dataset preparation, and experiments Curated from Kaggle competitions covering various domains (NLP, CV, signal processing), with human baselines from public leaderboards Challenging tasks testing real-world ML engineering skills; includes competitions like COVID-19 vaccine prediction and ancient scroll deciphering Competition-level; open-ended and difficult tasks for evaluating autonomous ML engineering
@@
PaperBench [313] 2025; 20 ICML 2024 papers, 8,316 individually gradable tasks Multi-language; AI research replication tasks Selected 20 Spotlight and Oral papers from ICML 2024, with rubrics co-developed with original paper authors Evaluates agents’ ability to replicate cutting-edge AI research from scratch, including understanding contributions, developing codebases, and executing experiments; uses LLM-based judge for automated grading Research-level; hierarchical rubrics decompose tasks into fine-grained sub-tasks
@@
Terminal-Bench [314] 2025; 80 tasks Multi-language; terminal environment tasks Hand-crafted and human-verified tasks, each with dedicated Docker environment, human-verified solutions, and test cases Evaluates agents on complete, end-to-end tasks in terminal environments including scientific workflows, network configuration, data analysis, API calls, and cybersecurity; emphasizes real-world complexity System-level; covers diverse terminal behaviors, directly executable in sandboxed environments
@@
SWE-RL [315] 2024; aggregated 24M PR instances and cloned 4.6M repositories Multi-language bug fixing tasks Crawled GitHub events from GHArchive (2015-2024), aggregated PRs and used heuristics to select bug-fixing seeds; each instance contains problem description, code context, and patches Uses difflib as reward function; used for training large models like Llama3-SWE-RL Project-level; massive scale
@@ 4.2.2 Continual Pre-training Strategies
Continual pre-training (CPT) enables models initially pre-trained on broad general corpora to undergo additional rounds of training on new, often domain-specific datasets, representing an important paradigm [316]. This approach enables models to acquire substantial new knowledge in specialized domains such as coding, mathematics, or scientific reasoning while preserving and building upon existing foundational capabilities [316], with research demonstrating that LLMs possess inherent abilities to effectively accumulate and retain knowledge across sequential training tasks [317, 1].
@@
The code domain is one of the most successful application areas for continual pre-training strategies. CodeLlama demonstrates this success by leveraging CPT on the general-purpose LLaMA2 foundation model to achieve state-of-the-art code generation performance [318, 249], while DeepSeek-Coder-V2 shows more advanced CPT capabilities by continuing training from DeepSeek-V2 intermediate checkpoints with an additional 6 trillion tokens, resulting in substantial enhancements to both coding abilities and mathematical reasoning capabilities [318, 249]. Critical to effective CPT are sophisticated data mixing strategies that carefully balance different data types and domains, with Qwen2.5-Coder’s systematic experimentation determining optimal 7:2:1 mixing ratios for code, text, and mathematics data, resulting in over 20% average performance improvement compared to code-only training approaches [318].
@@
However, preventing catastrophic forgetting of previously learned knowledge remains a central challenge in CPT [319, 320, 321]. This is addressed through strategies such as DeepSeek’s approach of replaying 30% of original pre-training data during continual training phases to maintain performance on previously mastered tasks [318], with replay-based methods being a well-established approach for mitigating catastrophic forgetting [322].
@@
Critically, data selection for continual pre-training leverages both distributed corpus features and linguistic characteristics to identify the most valuable training examples [323, 324, 325], with empirical research demonstrating that CPT on appropriately selected unlabeled task data and augmented unlabeled data sampled from in-domain corpora can significantly enhance end-task performance [323, 326]. Critical CPT challenges that remain active research areas include the "stability gap" phenomenon where models experience temporary but significant performance drops at the onset of new domain data training before eventually recovering and surpassing previous performance levels [327], the persistent issue of catastrophic forgetting of previous knowledge though research indicates self-supervised CPT approaches tend to outperform supervised protocols in knowledge retention [317, 1, 328], and the fundamental difficulty of scaling CPT methods to scenarios with small limited-token corpora where the restricted diversity of training examples significantly hinders effective knowledge acquisition and generalization [329, 330].
@@ 4.3.1 Supervised Fine-tuning
Large language model training follows well-established two-stage paradigms, with an initial pre-training phase on massive diverse corpora to acquire broad linguistic and world knowledge, followed by post-training that adapts models to excel at specific tasks and domains [58]. Supervised fine-tuning (SFT) is a critical technique that enhances LLM capabilities while aligning model behaviors with human expectations and requirements [225, 1, 59]. The SFT process involves adjusting pre-trained model parameters in a supervised manner by utilizing carefully curated datasets containing high-quality labeled examples, enabling models to maintain their broad foundational knowledge while simultaneously acquiring highly targeted specialization for specific applications [61, 60, 2, 331].
@@
Instruction tuning represents a particularly impactful form of supervised fine-tuning that enhances model capabilities by explicitly training models to understand and accurately follow diverse human instructions across varied task formats [279, 60], with research demonstrating substantial improvements in zero-shot performance on previously unseen task types [60, 59].
@@
Code generation is one of the most prominent and successful applications of supervised fine-tuning techniques, with notable implementations including CodeGen and StarCoder which employ a two-stage approach of initial pre-training on diverse multilingual code corpora followed by targeted fine-tuning on carefully selected monolingual data to enhance language-specific code generation quality [62], while systems like WizardCoder demonstrate alternative approaches by fine-tuning on specialized instruction corpora systematically derived from more capable teacher models [62, 2]. The extremely large parameter counts characteristic of modern language models make traditional approaches of updating all model parameters during fine-tuning computationally impractical and prohibitively expensive [63, 64], driving rapid development of Parameter-Efficient Fine-Tuning (PEFT) methodologies that achieve comparable performance with reduced computational requirements [332]. Adapter Tuning represents one influential PEFT approach by introducing small trainable adapter modules that add only a minimal number of new trainable parameters for each specific task while strategically keeping the vast majority of original pre-trained parameters completely fixed [332, 64], while Low-Rank Adaptation (LoRA) achieves even greater parameter efficiency by imposing carefully designed low-rank constraints on weight updates, enabling effective fine-tuning with orders of magnitude fewer trainable parameters [332].
@@
The critical importance of data quality over mere quantity has become a central guiding principle in supervised fine-tuning research and practice [333, 334, 65], with sophisticated data selection and curation strategies proving essential for optimal results. The AlpaGasus approach uses this method by systematically using ChatGPT to automatically identify and filter out low-quality training examples, achieving significant performance improvements using only 9,000 carefully selected high-quality samples compared to much larger but lower-quality datasets [66]. Interestingly, recent empirical research has revealed that supervised fine-tuning on high-quality demonstrations can in many cases result in superior performance compared to more complex preference fine-tuning approaches, particularly evident in challenging coding tasks where direct supervision appears especially effective [335].
@@ 4.3.2 Reinforcement Learning
Modern post-training approaches encompass two primary paradigms: supervised fine-tuning based on demonstration data, and preference fine-tuning methods utilizing techniques like RLHF and DPO [335, 67, 2, 68, 69].
@@
Among these paradigms, reinforcement learning has become a pivotal technique for post-training large language models, representing fundamental paradigm shifts away from traditional supervised fine-tuning approaches toward more dynamic optimization strategies [2, 70, 336, 71, 337]. While pre-training establishes robust foundational linguistic knowledge and pattern recognition capabilities, subsequent post-training refinement through sophisticated RL techniques has proven essential for substantially enhancing models’ complex reasoning abilities and adaptability to challenging real-world tasks [338]. The foundational framework for RL-based post-training was firmly established by RLHF systems that align model behaviors using carefully constructed reward signals derived from human preference judgments [71, 2, 339], with recent breakthrough systems like OpenAI’s o-series models and DeepSeek-R1 demonstrating that models trained with RL post-training can significantly outperform comparable models without RL training on demanding reasoning and coding benchmarks [71, 340].
@@
The reinforcement learning algorithmic landscape for language model training centers around several key complementary approaches, each with distinct advantages and trade-offs. PPO remains the most widely adopted baseline approach due to its carefully designed clipped objective functions that effectively stabilize the inherently unstable RL training process [71, 2, 341, 70], while DPO represents a major algorithmic simplification by cleverly replacing complex RL rollout procedures with more straightforward classification-style loss functions that directly optimize preference rankings [71, 68, 342]. Group Relative Policy Optimization (GRPO) emerged to address specific limitations in DPO by incorporating group-wise comparative feedback that enables more nuanced preference learning [71, 341]. Significant paradigm shifts have emerged with approaches like Reinforcement Learning with Verifiable Rewards (RLVR) that strategically replace learned reward models with deterministic task-specific rule-based reward functions, proving particularly effective for domains like coding and mathematics where answer correctness can be objectively verified [343].
@@
Reinforcement learning has become a transformative strategy for code generation applications, with significant empirical advancements demonstrated across multiple diverse programming domains and task types [336]. Foundational approaches in this area are exemplified by CodeRL, which employs sophisticated actor-critic framework architectures where critic networks learn to accurately predict functional correctness of generated code, providing learned reward signals that guide actor networks during the training process [341, 245]. Building on these foundations, PPOCoder demonstrates effective integration of Proximal Policy Optimization with direct execution-based feedback by utilizing actual compiler results and test case outcomes as immediate reward signals, enabling models to learn from real program behavior [341], while more recent developments have introduced sophisticated multi-stage training strategies exemplified by StepCoder which implements carefully designed curricula progressing through increasingly complex code completion subtasks [341]. Applications of RL-based code generation have extended to highly specialized technical domains, with systems like VeriSeek demonstrating RL’s effectiveness for generating correct hardware description language code in challenging Verilog generation tasks [344].
@@
RL-based post-training faces numerous fundamental challenges that remain active areas of research. Training stability issues include catastrophic forgetting of previously learned capabilities and reward hacking behaviors where models exploit reward function specifications rather than learning intended behaviors [338]. Scale-dependent effectiveness presents another significant limitation, with substantial benefits from RL training observed primarily in large models ranging from 8 billion to 670 billion parameters [345], while RL techniques remain considerably more challenging to apply successfully to smaller models with one billion or fewer parameters where training instability and sample efficiency become critical bottlenecks. Additionally, optimization dilemmas persist in the form of exploration-exploitation tradeoffs, with RL training often suffering from severe sample inefficiency particularly when initial policy distributions struggle to adequately explore regions of the action space containing high-reward trajectories [340], and fundamental mismatches between training objectives and evaluation metrics, with standard RL formulations focused on maximizing expected cumulative reward while practical evaluation of code generation systems typically relies on risk-seeking metrics like Pass@k that prioritize generating at least one correct solution among multiple attempts [339, 346, 347].

### v2

@@ 4.2.1 Pre-training Objectives
LeetCodeDataset [xia2025leetcodedatasettemporaldatasetrobust] 2025; 2,613 training problems and 256 test problems Python LeetCode algorithmic problems Wrapped LeetCode problems into data structures containing imports, entry_point, tests, completion, examples, and metadata Each problem provides complete unit tests and target code; suitable for model training and evaluation Function-level; includes Python test scripts
@@
OpenCodeReasoning [ahmad2025opencodereasoning] 2024; 735,255 samples covering 28,319 programming problems Multi-language programming problems with reasoning Integrated data from 11 programming platforms and generated solutions using NVIDIA R1 Large-scale high-quality programming reasoning data; supports training models with reasoning capabilities Problem-level; includes problem text, solutions, and reasoning
@@
DeepCoder [deepcoder2025] 2025; 24k verified coding problems, each with $\geq$5 test cases Competitive programming problems Merged TACO verified, Prime Intellect synthetic problems, LiveCodeBench, etc., filtered to 24k problems with complete tests Provides fully open high-quality data for RL training; reward uses binary model (pass/fail) Problem-level; includes complete test sets
@@
SWE-Bench Variants [jimenez2023swe, yang2024swebench_multimodal, yang2025swesmith] Since 2023; Full: 2,294, Lite: 300, Verified: 500, Multimodal: 617, Multilingual: 300 Primarily Python, also multi-language versions; project-level issue resolution Collected issues and PRs from GitHub, recording base commits, patches, and test patches; some instances expert-verified Instances include problem descriptions, base commits, patches, tests, and metadata; Multilingual version extends to 9 languages Project-level; used for evaluating bug-fixing capabilities of coding agents
@@
Multi-SWE-Bench [zan2025multi] 2025; 1,632 high-quality instances Seven languages: Java, TypeScript, JavaScript, Go, Rust, C, C++ Five-stage process: repository selection, PR crawling, environment determination, PR filtering, and human validation Fills the Python-only gap in SWE-Bench; cross-language evaluation for coding agents Project-level; used for evaluation
@@
MLE-bench [chan2024mle] 2024; 75 Kaggle competitions, 3.3TB dataset Multi-language; ML engineering tasks including model training, dataset preparation, and experiments Curated from Kaggle competitions covering various domains (NLP, CV, signal processing), with human baselines from public leaderboards Challenging tasks testing real-world ML engineering skills; includes competitions like COVID-19 vaccine prediction and ancient scroll deciphering Competition-level; open-ended and difficult tasks for evaluating autonomous ML engineering
@@
PaperBench [starace2025paperbench] 2025; 20 ICML 2024 papers, 8,316 individually gradable tasks Multi-language; AI research replication tasks Selected 20 Spotlight and Oral papers from ICML 2024, with rubrics co-developed with original paper authors Evaluates agents’ ability to replicate cutting-edge AI research from scratch, including understanding contributions, developing codebases, and executing experiments; uses LLM-based judge for automated grading Research-level; hierarchical rubrics decompose tasks into fine-grained sub-tasks
@@
Terminal-Bench [tbench2025] 2025; 80 tasks Multi-language; terminal environment tasks Hand-crafted and human-verified tasks, each with dedicated Docker environment, human-verified solutions, and test cases Evaluates agents on complete, end-to-end tasks in terminal environments including scientific workflows, network configuration, data analysis, API calls, and cybersecurity; emphasizes real-world complexity System-level; covers diverse terminal behaviors, directly executable in sandboxed environments
@@
SWE-RL [wei2025swe] 2024; aggregated 24M PR instances and cloned 4.6M repositories Multi-language bug fixing tasks Crawled GitHub events from GHArchive (2015-2024), aggregated PRs and used heuristics to select bug-fixing seeds; each instance contains problem description, code context, and patches Uses difflib as reward function; used for training large models like Llama3-SWE-RL Project-level; massive scale
@@ 4.2.2 Continual Pre-training Strategies
Continual pre-training (CPT) enables models initially pre-trained on broad general corpora to undergo additional rounds of training on new, often domain-specific datasets, representing an important paradigm [gao2024enhancing]. This approach enables models to acquire substantial new knowledge in specialized domains such as coding, mathematics, or scientific reasoning while preserving and building upon existing foundational capabilities [gao2024enhancing], with research demonstrating that LLMs possess inherent abilities to effectively accumulate and retain knowledge across sequential training tasks [singh2025beyond, brown2020language].
@@
The code domain is one of the most successful application areas for continual pre-training strategies. CodeLlama demonstrates this success by leveraging CPT on the general-purpose LLaMA2 foundation model to achieve state-of-the-art code generation performance [ibrahim2024simple, deepseekai2024deepseek], while DeepSeek-Coder-V2 shows more advanced CPT capabilities by continuing training from DeepSeek-V2 intermediate checkpoints with an additional 6 trillion tokens, resulting in substantial enhancements to both coding abilities and mathematical reasoning capabilities [ibrahim2024simple, deepseekai2024deepseek]. Critical to effective CPT are sophisticated data mixing strategies that carefully balance different data types and domains, with Qwen2.5-Coder’s systematic experimentation determining optimal 7:2:1 mixing ratios for code, text, and mathematics data, resulting in over 20% average performance improvement compared to code-only training approaches [ibrahim2024simple].
@@
However, preventing catastrophic forgetting of previously learned knowledge remains a central challenge in CPT [kirkpatrick2017overcoming, mccloskey1989catastrophic, van2024continual]. This is addressed through strategies such as DeepSeek’s approach of replaying 30% of original pre-training data during continual training phases to maintain performance on previously mastered tasks [ibrahim2024simple], with replay-based methods being a well-established approach for mitigating catastrophic forgetting [lesort2020continual].
@@
Critically, data selection for continual pre-training leverages both distributed corpus features and linguistic characteristics to identify the most valuable training examples [xie2023efficient, ruder2017learning, tsvetkov2016learning], with empirical research demonstrating that CPT on appropriately selected unlabeled task data and augmented unlabeled data sampled from in-domain corpora can significantly enhance end-task performance [xie2023efficient, gururangan2020don]. Critical CPT challenges that remain active research areas include the "stability gap" phenomenon where models experience temporary but significant performance drops at the onset of new domain data training before eventually recovering and surpassing previous performance levels [guo2024efficient], the persistent issue of catastrophic forgetting of previous knowledge though research indicates self-supervised CPT approaches tend to outperform supervised protocols in knowledge retention [singh2025beyond, brown2020language, mirzadeh2021wide], and the fundamental difficulty of scaling CPT methods to scenarios with small limited-token corpora where the restricted diversity of training examples significantly hinders effective knowledge acquisition and generalization [ovadia2025knowledge, kandpal2022large].
@@ 4.3.1 Supervised Fine-tuning
Large language model training follows well-established two-stage paradigms, with an initial pre-training phase on massive diverse corpora to acquire broad linguistic and world knowledge, followed by post-training that adapts models to excel at specific tasks and domains [wei2025vericoder]. Supervised fine-tuning (SFT) is a critical technique that enhances LLM capabilities while aligning model behaviors with human expectations and requirements [shengyu2023instruction, brown2020language, chung2022scaling]. The SFT process involves adjusting pre-trained model parameters in a supervised manner by utilizing carefully curated datasets containing high-quality labeled examples, enabling models to maintain their broad foundational knowledge while simultaneously acquiring highly targeted specialization for specific applications [wu2024sparkra, wei2021finetuned, ouyang2022training, liu2024m2rc].
@@
Instruction tuning represents a particularly impactful form of supervised fine-tuning that enhances model capabilities by explicitly training models to understand and accurately follow diverse human instructions across varied task formats [jiang2024survey, wei2021finetuned], with research demonstrating substantial improvements in zero-shot performance on previously unseen task types [wei2021finetuned, chung2022scaling].
@@
Code generation is one of the most prominent and successful applications of supervised fine-tuning techniques, with notable implementations including CodeGen and StarCoder which employ a two-stage approach of initial pre-training on diverse multilingual code corpora followed by targeted fine-tuning on carefully selected monolingual data to enhance language-specific code generation quality [shen2023pangu], while systems like WizardCoder demonstrate alternative approaches by fine-tuning on specialized instruction corpora systematically derived from more capable teacher models [shen2023pangu, ouyang2022training]. The extremely large parameter counts characteristic of modern language models make traditional approaches of updating all model parameters during fine-tuning computationally impractical and prohibitively expensive [hu2022lora, houlsby2019parameter], driving rapid development of Parameter-Efficient Fine-Tuning (PEFT) methodologies that achieve comparable performance with reduced computational requirements [yuan2023evaluating]. Adapter Tuning represents one influential PEFT approach by introducing small trainable adapter modules that add only a minimal number of new trainable parameters for each specific task while strategically keeping the vast majority of original pre-trained parameters completely fixed [yuan2023evaluating, houlsby2019parameter], while Low-Rank Adaptation (LoRA) achieves even greater parameter efficiency by imposing carefully designed low-rank constraints on weight updates, enabling effective fine-tuning with orders of magnitude fewer trainable parameters [yuan2023evaluating].
@@
The critical importance of data quality over mere quantity has become a central guiding principle in supervised fine-tuning research and practice [nakkiran2019deep, shumailov2024models, zhou2023lima], with sophisticated data selection and curation strategies proving essential for optimal results. The AlpaGasus approach uses this method by systematically using ChatGPT to automatically identify and filter out low-quality training examples, achieving significant performance improvements using only 9,000 carefully selected high-quality samples compared to much larger but lower-quality datasets [chen2023alpagasus]. Interestingly, recent empirical research has revealed that supervised fine-tuning on high-quality demonstrations can in many cases result in superior performance compared to more complex preference fine-tuning approaches, particularly evident in challenging coding tasks where direct supervision appears especially effective [xiao2024how].
@@ 4.3.2 Reinforcement Learning
Modern post-training approaches encompass two primary paradigms: supervised fine-tuning based on demonstration data, and preference fine-tuning methods utilizing techniques like RLHF and DPO [xiao2024how, iacovides2025findpo, ouyang2022training, rafailov2023direct, xiong2025rag].
@@
Among these paradigms, reinforcement learning has become a pivotal technique for post-training large language models, representing fundamental paradigm shifts away from traditional supervised fine-tuning approaches toward more dynamic optimization strategies [ouyang2022training, stiennon2020learning, wang2024enhancing, wang2025dump, dou2025improving]. While pre-training establishes robust foundational linguistic knowledge and pattern recognition capabilities, subsequent post-training refinement through sophisticated RL techniques has proven essential for substantially enhancing models’ complex reasoning abilities and adaptability to challenging real-world tasks [ferrag2025reasoning]. The foundational framework for RL-based post-training was firmly established by RLHF systems that align model behaviors using carefully constructed reward signals derived from human preference judgments [wang2025dump, ouyang2022training, xi2025agentgym], with recent breakthrough systems like OpenAI’s o-series models and DeepSeek-R1 demonstrating that models trained with RL post-training can significantly outperform comparable models without RL training on demanding reasoning and coding benchmarks [wang2025dump, xu2025kdrl].
@@
The reinforcement learning algorithmic landscape for language model training centers around several key complementary approaches, each with distinct advantages and trade-offs. PPO remains the most widely adopted baseline approach due to its carefully designed clipped objective functions that effectively stabilize the inherently unstable RL training process [wang2025dump, ouyang2022training, rodriguez2025rendering, stiennon2020learning], while DPO represents a major algorithmic simplification by cleverly replacing complex RL rollout procedures with more straightforward classification-style loss functions that directly optimize preference rankings [wang2025dump, rafailov2023direct, gao2025one]. Group Relative Policy Optimization (GRPO) emerged to address specific limitations in DPO by incorporating group-wise comparative feedback that enables more nuanced preference learning [wang2025dump, rodriguez2025rendering]. Significant paradigm shifts have emerged with approaches like Reinforcement Learning with Verifiable Rewards (RLVR) that strategically replace learned reward models with deterministic task-specific rule-based reward functions, proving particularly effective for domains like coding and mathematics where answer correctness can be objectively verified [taveekitworachai2025prior].
@@
Reinforcement learning has become a transformative strategy for code generation applications, with significant empirical advancements demonstrated across multiple diverse programming domains and task types [wang2024enhancing]. Foundational approaches in this area are exemplified by CodeRL, which employs sophisticated actor-critic framework architectures where critic networks learn to accurately predict functional correctness of generated code, providing learned reward signals that guide actor networks during the training process [rodriguez2025rendering, le2022coderl]. Building on these foundations, PPOCoder demonstrates effective integration of Proximal Policy Optimization with direct execution-based feedback by utilizing actual compiler results and test case outcomes as immediate reward signals, enabling models to learn from real program behavior [rodriguez2025rendering], while more recent developments have introduced sophisticated multi-stage training strategies exemplified by StepCoder which implements carefully designed curricula progressing through increasingly complex code completion subtasks [rodriguez2025rendering]. Applications of RL-based code generation have extended to highly specialized technical domains, with systems like VeriSeek demonstrating RL’s effectiveness for generating correct hardware description language code in challenging Verilog generation tasks [wang2024large].
@@
RL-based post-training faces numerous fundamental challenges that remain active areas of research. Training stability issues include catastrophic forgetting of previously learned capabilities and reward hacking behaviors where models exploit reward function specifications rather than learning intended behaviors [ferrag2025reasoning]. Scale-dependent effectiveness presents another significant limitation, with substantial benefits from RL training observed primarily in large models ranging from 8 billion to 670 billion parameters [le2025reasoning], while RL techniques remain considerably more challenging to apply successfully to smaller models with one billion or fewer parameters where training instability and sample efficiency become critical bottlenecks. Additionally, optimization dilemmas persist in the form of exploration-exploitation tradeoffs, with RL training often suffering from severe sample inefficiency particularly when initial policy distributions struggle to adequately explore regions of the action space containing high-reward trajectories [xu2025kdrl], and fundamental mismatches between training objectives and evaluation metrics, with standard RL formulations focused on maximizing expected cumulative reward while practical evaluation of code generation systems typically relies on risk-seeking metrics like Pass@k that prioritize generating at least one correct solution among multiple attempts [xi2025agentgym, qi2024webrl, xia2024inverse].